All Seasons Plumbing has two service trucks that frequently need repair. If the probability the first truck is available is 0.66, the probability the second truck is available is 0.45, and the probability that both trucks are available is 0.37:

What is the probability neither truck is available? (Round your answer to 2 decimal places.)

Answer by stanbon(75887)

All Seasons Plumbing has two service trucks that frequently need repair. If the probability the first truck is available is 0.66, the probability the second truck is available is 0.45, and the probability that both trucks are available is 0.37:
What is the probability neither truck is available? (Round your answer to 2 decimal places.)
-----
P(1styes AND 2ndno) = 0.66*0.55 = 0.36
----
P(1styes AND 2ndyes) = 0.37
----
P(1stno AND 2ndyes) = 0.34*0.45 = 0.15
------
Ans: P(1stno AND 2ndno) = 1-[0.36+0.37+0.15] = 0.12
--------------------------------
Cheers,
Stan H.